4x+3y=-5
-9x+y =50
solve using the substitution method

Answer by rfer(16322)

-9x+y=50 y=50+9x
4x+3(50+9x)=-5
4x+150+27x=-5
31x=-155
x=-5
---------------
-9*-5+y=50
45+y=50
y=5
---------------
proof
-9*-5+5=50